# Coreset Strikes Back: Improved Parameterized Approximation Schemes for (Constrained) $k$-Median/Means

### v2

@@ 3.3 Analysis
In this section, we analyze the running time and correctness of Algorithm 1. First, in Section 3.3.1, we show that Algorithm 1 has bounded running time (Lemma 11), and later, in Section 3.3.2, we show that there there exists an iteration of for loop in 4 of Algorithm 1 that does not return $\bot$ (Lemma 12). Finally, we use these results to prove Theorem 10.
For keeping the analysis clean, we show the proof for Well Constrained $(k,z)$-Median. We remark that the same analysis can be extended easily to Well Constrained $(k,z)$-Clustering, by replacing $\varepsilon$ with $\varepsilon/z$. Let $\mathcal{I}$ be an instance of Well Constrained $(k,z)$-Clustering that is given as an input to Algorithm 1, and let $(Y,w)$ be the corresponding coreset for $\mathcal{I}$ obtained in Step 3.
@@ 3.3.1 Running time
[…]
@@
[…]
@@
[…]
@@ 3.3.2 Correctness
Let $\mathcal{O}=(O=(o_{1},\dots,o_{k}),f^{*})$ be an optimal solution to coreset $(Y,w)$ with cost ${\sf OPT}$. For $i\in[k]$, let $p^{*}_{i}=\arg\min_{p\in P}\{d(p,o_{i})\mid f^{*}\text{ assigns $p$ to $o_{i}$}\}$, be the leader of (the cluster corresponding to) $o_{i}$, by breaking ties arbitrarily. Similarly, let $r_{i}=d(p^{*}_{i},o_{i})$ be the distance of $o_{i}$ to its leader $p^{*}_{i}$. Now, we fix the iteration of the for loop of Algorithm 1 corresponding to $(p^{\prime}_{1},\dots,p^{\prime}_{k})=(p^{*}_{1},\dots,p^{*}_{k})$ and $(r^{\prime}_{1},\dots,r^{\prime}_{k})=(\tilde{r}_{1},\dots,\tilde{r}_{k})\in[\Delta]^{k}_{\varepsilon}$, where $r_{i}\leq\tilde{r}_{i}<(1+\varepsilon)r_{i},i\in[k]$. Let $\mathcal{T}$ be the recursion tree of COMPUTE-SOLUTION corresponding to this iteration with root $t_{0}$.
Observation 13.
@@
If ${\sf cost}(Y,X,f)>(1+5\varepsilon)\cdot\mathcal{G}$, then there exists a point $p\in Y$ and an index $i\in[k]$ satisfying the following properties.
@@
Let $W\coloneqq\left\{(p,i)\in Y\times[k]:\text{$f^{*}$ assigns $p$ to $i$, and }{\sf d}(p,x_{i})>(1+\varepsilon)\cdot{\sf d}(p,o_{i})\right\}$. First, we show that $W$ is non-empty. Suppose for the sake of contradiction that $U$ is empty, i.e., for each $p\in Y$ and all $i\in[k]$ such that $f^{*}$ assigns $p$ to $o_{i}$, it holds that ${\sf d}(p,x_{i})\leq(1+\varepsilon)\cdot{\sf d}(p,o_{i})$. Then, by the properties of BALL-INT and ASSIGNMENT algorithms, it follows that $(X,f^{*})$ is feasible. Hence, ${\sf cost}(Y,X,f^{*})\leq(1+\varepsilon){\sf cost}(Y,O,f^{*})$. Therefore, we have
${\sf cost}(Y,X,f)$ $\leq(1+\varepsilon)\cdot{\sf cost}(Y,X,f^{*})$ due to ASSIGNMENT
$\leq(1+\varepsilon)^{2}\cdot{\sf cost}(Y,O,f^{*})$
$\leq(1+\varepsilon)^{3}\cdot{\sf cost}(P,O,f^{*})$ due to coreset
$\leq(1+\varepsilon)^{3}\cdot\mathcal{G}$ since $\text{since }\mathcal{G}\geq{\sf OPT}$
$\leq(1+5\varepsilon)\cdot\mathcal{G}$ since $\text{since }\varepsilon\leq 1/2$
@@
as desired, since $r_{i}\leq\tilde{r}_{i}<(1+\varepsilon)r_{i}$ due to 13. Now, suppose the property $(d)$ does not hold, i.e., ${\sf d}(p,p^{*}_{i})>\frac{6\tilde{r}_{i}}{\varepsilon}$. Then, we obtain the following by triangle inequality.
@@
Fix some $(p_{t},i_{t})\in Y\times[k]$ satisfying 14.1 and note that $p_{t}\in B_{i_{t}}\setminus{\sf ball}(x_{i_{t}},\tilde{r}_{i_{t}})$. Hence, consider the corresponding iteration of the for loop to $(p_{t},i_{t})$ in 5, and let $t^{\prime}\in\mathcal{T}$ be the corresponding child of $t$ in this iteration. We claim that $t^{\prime}$ is consistent with $O$. Towards this, let $(Q_{1},\dots,Q_{k})$ and $(Q^{\prime}_{1},\dots,Q^{\prime}_{k})$ be first arguments at $t$ and $t^{\prime}$, respectively. Then, note that for $i\neq i_{t}$, we have, $Q^{\prime}_{i}=Q_{i}$, and $Q^{\prime}_{i_{t}}=Q_{i_{t}}\cup\{(p_{t},\alpha_{t})\}$, where $\alpha_{t}=\frac{{\sf d}(p_{t},x_{i_{t}})}{1+\varepsilon}$. Hence, for any $i\neq i_{t}$, we have that for $(p,\alpha)\in Q^{\prime}_{i}=Q_{i}$, it holds that ${\sf d}(p,o_{i})\leq\alpha$ due to the induction hypothesis at $t$. Similarly, for $i_{t}$, we have for any $(p,\alpha)\in Q^{\prime}_{i_{t}}\setminus\{(p_{t},\alpha_{t})\}$, it holds that ${\sf d}(p,o_{i})\leq\alpha$ due to the induction hypothesis at $t$. Now, consider $(p_{t},\alpha_{t})\in Q^{\prime}_{i_{t}}$, and note that $\alpha_{t}=\frac{{\sf d}(p_{t},x_{i_{t}})}{1+\varepsilon}>{\sf d}(p_{t},o_{i_{t}})$ since ${\sf d}(p_{t},x_{i_{t}})>(1+\varepsilon)\cdot{\sf d}(p_{t},o_{i_{t}})$ from 14.1, finishing the proof of the claim. Finally, note that since none of the nodes of path $\mathcal{P}$ return $\bot$, the root node $t_{0}$ never returns $\bot$, finishing the proof of the lemma.
@@
Suppose Algorithm 1 terminates with a solution $(S,f_{S})$. Then, we have that ${\sf cost}(Y,S)\leq(1+5\varepsilon)\cdot\mathcal{G}\leq(1+5\varepsilon)(1+\varepsilon){\sf OPT}$. Therefore, we have
${\sf cost}(P,S,f_{S})$ $\leq(1+\varepsilon){\sf cost}(P,S)$ due to ASSIGNMENT
@@ 4 Applications of the Framework
In this section, we discuss some applications of the framework.
4.1 Capacitated Clustering
Uniform Capacities.
@@ 4.1 Capacitated Clustering
This is the most straightforward application of our framework. For capacitated $k$-median, a coreset of size $\left(\frac{k\log n}{\varepsilon}\right)^{O(1)}$ in [CL19a]. Note that this coreset is for general metric spaces, and even works with non-uniform capacities. Further, it is possible to generalize these coresets for $(k,z)$-clustering for any fixed $z\geq 1$, where the size of the coreset then depends on $z$. The Assignment subroutine in this case is simply the minimum-cost maximum flow problem, which is solvable in polynomial time, and produces integral flows. Finally, we note that uniform capacities is also the only sensible notion in continuous versions of capacitated $k$-median/means, where the set of candidate facilities is infinite, e.g., $\mathbb{R}^{d}$.
Non-Uniform Capacities.
Here, we need to slightly adapt the framework, for which we follow the approach taken in [CL19a]. We start with the coreset of size $\left(\frac{k\log n}{\varepsilon}\right)^{O(1)}$ from the same paper. Then, we use the technique of color coding, as follows. For each $c\in F$, we assign a color $i\in[k]$ chosen uniformly at random. A good coloring corresponds to a coloring in which all centers of a fixed optimal solution $O\subseteq F$ of size $k$ receieve distinct colors. Note that this happens with probability $1/e^{k}$, and this procedure can be derandomized by using standard tools from parameterized complexity [CFK+15]. Thus, now we assume that we have a good coloring, and we want to find a colorful set of $k$ centers. Then, in the algorithm, we make the following change. BALL-INT algorithm for the $i$th request set $Q_{i}$, selects a center that (i) satisfies all the requests in $Q_{i}$, and (ii) has a maximum capacity among all centers with color $i$. This ensures that, assuming all the guesses were correct, for each $i\in[k]$, we select a center $x_{i}$ such that (i) for each $p\in Y$, if $f^{*}$ assigns $p$ to $i$, then ${\sf d}(p,x_{i})\leq(1+\varepsilon)\cdot{\sf d}(p,o_{i})$, and (ii) capacity of $x_{i}$ is at least that of $o_{i}$. This guarantees that there is a feasible assignment from $Y$ to $X$ of cost at most $(1+O(\varepsilon))$ times ${\sf OPT}$.
4.2 $(\alpha,\beta)$-Fair Clustering and Variants

### v3

@@ 3.3 Analysis
In this section, we analyze the running time and correctness of Algorithm 1. First, in Section 3.3.1, we show that Algorithm 1 has bounded running time (Lemma 12), and later, in Section 3.3.2, we show that there exists an iteration of for loop in 4 of Algorithm 1 that does not return $\bot$ (Lemma 13). Finally, we use these results to prove Theorem 11.
For keeping the analysis clean, we show the proof for Well Constrained $(k,z)$-Median. We remark that the same analysis can be extended easily to Well Constrained $(k,z)$-Clustering, by replacing $\varepsilon$ with $\varepsilon/z$. Let $\mathcal{I}$ be an instance of Well Constrained $(k,z)$-Clustering that is given as an input to Algorithm 1, and let $(Y,w)$ be the corresponding coreset for $\mathcal{I}$ obtained in 3.
@@ 3.3.1 Running time
For an execution of $\texttt{COMPUTE-SOLUTION}((Q_{1},\dots,Q_{k}),\mathcal{B})$, we call the first argument $(Q_{1},\dots,Q_{k})$ of the execution as the request set of the execution. Towards this, we show the following claim.
@@
Consider the execution of an iteration of the for loop of Algorithm 1 with tuple $(p^{\prime}_{1},\dots,p^{\prime}_{k})$ and distances $(r^{\prime}_{1},\dots,r^{\prime}_{k})$ and the corresponding invocation of COMPUTE-SOLUTION with arguments $(Q_{1},\dots,Q_{k})$ and $\mathcal{B}$. Let $\mathcal{T}$ be the corresponding recursion tree with root node $t_{0}$. For any node $t\in\mathcal{T}$, we denote by $(Q^{t}_{1},\cdots,Q^{t}_{k})$ as the request set with which it is invoked during the execution. Furthermore, we call the request set corresponding to root node $t_{0}$ as the initial request set of $\mathcal{T}$. Notice that the request set of any non-root node $t\in\mathcal{T}$ differs with the request set of its parent node $t^{\prime}$ in exactly one request. More specifically, there exists $j\in[k]$ such that $Q^{t}_{j}=Q^{t^{\prime}}_{j}\cup\{(p,\alpha)\}$, for some $(p,\alpha)\in Y\times[\Delta]_{\varepsilon}$, and $Q^{t}_{i}=Q^{t^{\prime}}_{i}$, for $i\neq j$. In this case, we say that $t$ has a modified $Q_{j}$, and that $(p,\alpha)$ as the modifier of $t$. Consider a path $\mathcal{P}$ from the root of $\mathcal{T}$, and for $i\in[k]$, let $T_{i}=\{t^{1}_{i},t^{2}_{i},\dots\}\subseteq\mathcal{P}$ be the nodes in $\mathcal{P}$ that have modified $Q_{i}$. Let $x^{1}_{i},x^{2}_{i},\dots$ be the centers computed for cluster $i$ at nodes $t^{1}_{i},t^{2}_{i},\dots$, respectively (see 1), and let $(p^{1}_{i},\alpha^{1}_{i}),(p_{i}^{2},\alpha^{2}_{i}),\dots$ be the modifiers of nodes $t^{1}_{i},t^{2}_{i},\dots$, respectively (see 6). That is, $x^{j}_{i}$ is the center computed at $t^{j}_{i}\in\mathcal{P}$ for cluster $i$. Since, for $j>1$, node $t^{j}_{i}$ computes $X^{j}=(x^{j}_{1},\dots,x^{j}_{k})$ using GenBALL-INT such that $x^{j}_{i}$ satisfies the cluster constraint containing $(p^{1}_{i},\alpha^{1}_{i}),\dots,(p_{i}^{j-1},\alpha^{j-1}_{i})$ requests, i.e., it holds that ${\sf d}(x^{j}_{i},p^{j^{\prime}}_{i})\leq(1+\varepsilon)\alpha^{j^{\prime}}_{i}$, for $j^{\prime}<j$. Furthermore, since node $t^{j}_{i}$ adds $\left(p^{j}_{i},\frac{{\sf d}(p^{j}_{i},x^{j}_{i})}{(1+\varepsilon)}\right)$ to $Q_{i}$, it holds that $\alpha^{j}_{i}=\frac{{\sf d}(p^{j}_{i},x^{j}_{i})}{(1+\varepsilon)}$. Hence, ${\sf d}(p^{j}_{i},x^{j}_{i})>(1+\tfrac{\varepsilon}{2})\alpha^{j}_{i}$. Therefore, the sequence $(x^{1}_{i},p^{1}_{i},\alpha^{1}_{i}),(x^{2}_{i},p^{2}_{i},\alpha^{2}_{i}),\dots$ is a $(\texttt{BALL-INT}_{i},\tfrac{\varepsilon}{2})$-algorithmic scattering. Finally, since $p^{j}_{i}\in{\sf ball}(p^{\prime}_{i},\nicefrac{{6r^{\prime}_{i}}}{{\varepsilon}})\setminus{\sf ball}(x_{i},r^{\prime}_{i})$, where $(p^{\prime}_{i},r^{\prime}_{i})=Q^{t_{0}}_{i}$ is the $i^{th}$ request in the initial request set of $\mathcal{T}$, we have that, $\alpha^{j}_{i}=\frac{d(p^{j}_{i},x^{j}_{i})}{1+\varepsilon}\geq r^{\prime}_{i}$. On the other hand,
@@
where the last inequality follows since $d(p^{\prime}_{i},x^{j}_{i})\leq(1+\varepsilon)r^{\prime}_{i}$, due to the initial request $(p^{\prime}_{i},r^{\prime}_{i})\in Q^{t_{0}}_{i}$. Thus, the radii in the requests $(p^{1}_{i},\alpha^{1}_{i}),(p_{i}^{2},\alpha^{2}_{i}),\dots$ lie in the interval $\left[{r^{\prime}_{i}},{7r^{\prime}_{i}}/{\varepsilon}\right]$. Hence, using Lemma 10, the length of $(\texttt{BALL-INT}_{i},\varepsilon/2)$-algorithmic scattering $(x^{1}_{i},p^{1}_{i},\alpha^{1}_{i}),(x^{2}_{i},p^{2}_{i},\alpha^{2}_{i}),\dots$ is bounded by $O(\lambda(\tfrac{\varepsilon}{2})\tfrac{\log(1/\varepsilon)}{\varepsilon})$, since the $(\texttt{BALL-INT}_{i},\varepsilon/2)$-algorithmic dimension of $\mathcal{M}$ is bounded by $\lambda(\varepsilon)$ due to Lemma 9. Since every node in $\mathcal{P}$ has a modified $Q_{j},j\in[k]$, we have that the length of the path $\mathcal{P}$ is bounded by $O(k\lambda(\tfrac{\varepsilon}{2})\tfrac{\log(1/\varepsilon)}{\varepsilon})$. ∎
@@ 3.3.2 Correctness
For any feasible solution $(X,f)$, we say that $f$ assigns $p\in Y$ to $x\in X$ if $f(p,x)>0$, and denote $f(p,\cdot):=\{x\in X|f(p,x)>0\}$, the set of centers assigned to $p$ in $X$. Let $\mathcal{O}=(O=(o_{1},\dots,o_{k}),f^{*})$ be an optimal solution to coreset $(Y,w)$ with cost $\sf OPT$} [^17] For brevity, we abuse OPT to denote the cost of an optimal solution in coreset rather than that of the original instance. This is fine, since the cost of optimal solution in coreset is same as that in the input instance, up to a multiplicative factor of $(1+\varepsilon)$.. For $i\in[k]$, let $p^{*}_{i}=\arg\min_{p\in P}\{d(p,o_{i})\mid f^{*}\text{ assigns $p$ to $o_{i}$}\}$, be the leader of (the cluster corresponding to) $o_{i}$, by breaking ties arbitrarily. Similarly, let $r_{i}=d(p^{*}_{i},o_{i})$ be the distance of $o_{i}$ to its leader $p^{*}_{i}$. Now, we fix the iteration of the for loop of Algorithm 1 corresponding to $(p^{\prime}_{1},\dots,p^{\prime}_{k})=(p^{*}_{1},\dots,p^{*}_{k})$ and $(r^{\prime}_{1},\dots,r^{\prime}_{k})=(\tilde{r}_{1},\dots,\tilde{r}_{k})\in[\Delta]^{k}_{\varepsilon}$, where $r_{i}\leq\tilde{r}_{i}<(1+\varepsilon)r_{i},i\in[k]$. Let $\mathcal{T}$ be the recursion tree of COMPUTE-SOLUTION corresponding to this iteration with root $t_{0}$.
Observation 14.
@@
If ${\sf wcost}(Y,X,f)>(1+5\varepsilon)\cdot\mathcal{G}$, then there exists a point $p\in Y$ and an index $i\in[k]$ satisfying the following properties.
@@
Let $W\coloneqq\{(p,i)\in Y\times[k]:i\in f^{*}(p,\cdot)\text{ and }{\sf d}(p,x_{i})>(1+\varepsilon)\cdot{\sf d}(p,o_{i})\}$. First, we show that $W$ is non-empty. Suppose for the sake of contradiction that $W$ is empty. Since, $\sum_{j\in[k]}f^{*}(p,o_{j})=\sum_{j\in[k]}f^{*}(p,x_{j})=w(p)$, this means that, for each $p\in Y$ and all $i\in[k]$ such that $f^{*}$ assigns $p$ to $o_{i}$, it holds that ${\sf d}(p,x_{i})\leq(1+\varepsilon)\cdot{\sf d}(p,o_{i})$. Then, by the properties of GenBALL-INT and ASSIGNMENT algorithms, and due to their compatibility, it follows that $(X,f^{*})$ is a feasible solution. Furthermore, for all $p\in Y$, we have,
${\sf wcost}(p,X,f^{*})$ $=\sum_{i\in f^{*}(p,\cdot)}f^{*}(p,x_{i})\cdot d(p,i)$
$\leq\sum_{i\in f^{*}(p,\cdot)}(1+\varepsilon)\cdot d(p,o_{i})$
$=(1+\varepsilon)\cdot{\sf wcost}(p,O,f^{*})$
This means, ${\sf wcost}(Y,X,f^{*})=\sum_{p\in Y}{\sf wcost}(p,X,f^{*})\leq(1+\varepsilon)\sum_{p\in Y}{\sf wcost}(p,O,f^{*})=(1+\varepsilon)\cdot{\sf wcost}(Y,O,f^{*})$. Therefore, we have
${\sf wcost}(Y,X,f)$ $\leq(1+\varepsilon)\cdot{\sf cost}(Y,X,f^{*})$ due to ASSIGNMENT
$\leq(1+\varepsilon)^{2}\cdot{\sf wcost}(Y,O,f^{*})$
$\leq(1+\varepsilon)^{3}\cdot{\sf wcost}(P,O,f^{*})$ due to coreset
$\leq(1+\varepsilon)^{3}\cdot\mathcal{G}$ $\text{since }\mathcal{G}\geq{\sf OPT}$
$\leq(1+5\varepsilon)\cdot\mathcal{G}$ $\text{since }\varepsilon\leq 1/2$
@@
as desired, since $r_{i}\leq\tilde{r}_{i}<(1+\varepsilon)r_{i}$ due to 14. Now, suppose the property $(d)$ does not hold, i.e., ${\sf d}(p,p^{*}_{i})>\frac{6\tilde{r}_{i}}{\varepsilon}$. Then, we obtain the following by triangle inequality.
@@
Fix some $(p_{t},i_{t})\in Y\times[k]$ satisfying 15.1 and note that $p_{t}\in B_{i_{t}}\setminus{\sf ball}(x_{i_{t}},\tilde{r}_{i_{t}})$. Hence, consider the corresponding iteration of the for loop to $(p_{t},i_{t})$ in 5, and let $t^{\prime}\in\mathcal{T}$ be the corresponding child of $t$ in this iteration. We claim that $t^{\prime}$ is consistent with $O$. Towards this, let $(Q_{1},\dots,Q_{k})$ and $(Q^{\prime}_{1},\dots,Q^{\prime}_{k})$ be first arguments at $t$ and $t^{\prime}$, respectively. Then, note that for $i\neq i_{t}$, we have, $Q^{\prime}_{i}=Q_{i}$, and $Q^{\prime}_{i_{t}}=Q_{i_{t}}\cup\{(p_{t},\alpha_{t})\}$, where $\alpha_{t}=\frac{{\sf d}(p_{t},x_{i_{t}})}{1+\varepsilon}$. Hence, for any $i\neq i_{t}$, we have that for $(p,\alpha)\in Q^{\prime}_{i}=Q_{i}$, it holds that ${\sf d}(p,o_{i})\leq\alpha$ due to the induction hypothesis at $t$. Similarly, for $i_{t}$, we have for any $(p,\alpha)\in Q^{\prime}_{i_{t}}\setminus\{(p_{t},\alpha_{t})\}$, it holds that ${\sf d}(p,o_{i})\leq\alpha$ due to the induction hypothesis at $t$. Now, consider $(p_{t},\alpha_{t})\in Q^{\prime}_{i_{t}}$, and note that $\alpha_{t}=\frac{{\sf d}(p_{t},x_{i_{t}})}{1+\varepsilon}>{\sf d}(p_{t},o_{i_{t}})$ since ${\sf d}(p_{t},x_{i_{t}})>(1+\varepsilon)\cdot{\sf d}(p_{t},o_{i_{t}})$ from 15.1, finishing the proof of the claim. Finally, note that since none of the nodes of path $\mathcal{P}$ return $\bot$, the root node $t_{0}$ never returns $\bot$, finishing the proof of the lemma.
@@
Suppose Algorithm 1 terminates with a solution $(S,f_{S})$. Then, we have that ${\sf wcost}(Y,S)\leq(1+5\varepsilon)\cdot\mathcal{G}\leq(1+5\varepsilon)(1+\varepsilon){\sf OPT}$. Therefore, we have
${\sf wcost}(P,S,f_{S})$ $\leq(1+\varepsilon){\sf cost}(P,S)$ due to ASSIGNMENT
@@ 4 Applications of the Framework
In this section, we discuss some applications of the framework. Key constructions in these applications are designing GenBALL-INT and ASSIGNMENT subroutines that satisfy the requirements of our framework (Section 3.1). We divide our applications based on constraints on center selection.
4.1 Constrained Center Selection
In these problems, there are hard constraints on choosing centers as a solution. This forces designing GenBALL-INT (and BALL-INT) algorithm slightly challenging compared to vanilla clustering. For example,in the framework of [ABB+23], it is absolutely fine for the ball intersection algorithm to chose the same center for different request sets during the execution, since the centers do not have any constraints. However, this is not allowed in problems such as capacitated clustering and fault tolerant clustering, and hence need to be addressed, depending on the problems.
Preprocessing.
Towards this, we first transform the input instance $\mathcal{I}$ in metric space $M=(P,F,d)$ to another instance $\mathcal{I}^{\prime}$ in metric space $M^{\prime}=(P,F^{\prime},d)$ such that the facility set partitioned into $k$ sets, i.e., $F^{\prime}=\dot{\bigcup}_{t\in[k]}F^{\prime}_{t}$, and ask for a solution $S$ to $\mathcal{I}$ to satisfy $|S\cap F_{t}|=1$, for all $t\in[k]$, in addition to the existing constraints of $\mathcal{I}$. Finally, we want that the optimal cost of $\mathcal{I}^{\prime}$ is equal to that of $\mathcal{I}$. This can be achieved using color coding technique, as follows. For each $c\in F$, we assign a color $i\in[k]$ chosen uniformly at random. A good coloring corresponds to a coloring in which all centers of a fixed optimal solution $O\subseteq F$ of size $k$ receive distinct colors. Note that this happens with probability $1/e^{k}$, and this procedure can be derandomized in time $2^{O(k)}n\log n$ by using standard tools from parameterized complexity [CFK+15]. Thus, now we assume that by inuring an additive factor of $2^{O(k)}n\log n$ and a multiplicative factor of $2^{O(k)}\log n$ in time, the input instance has facility set partitioned into $k$ parts such that $O$ is colorful, and we want to find a colorful set of $k$ centers that is feasible for the given constraints.
4.1.1 Capacitated Clustering
@@
For capacitated $k$-median, a coreset of size $\left(\frac{k\log n}{\varepsilon}\right)^{O(1)}$ is known [CL19]. Note that this coreset is for general metric spaces, and even works with non-uniform capacities. Further, it is possible to generalize these coresets for $(k,z)$-clustering for any fixed $z\geq 1$, where the size of the coreset depends on $z$.
The Assignment subroutine in this case is simply the minimum-cost maximum flow problem, which is solvable in polynomial time, and produces integral flows.
The GenBALL-INT algorithm is composed of $k$ ball intersection algorithms $\{\texttt{BALL-INT}_{i}\}_{i\in[k]}$, one for every request set in $\{Q_{i}\}_{i\in[k]}$. $\texttt{BALL-INT}_{i},i\in[k]$ on request set $Q_{i}$, selects a center in $F_{i}$ with maximum capacity that satisfies all the requests in $Q_{i}$. This ensures that, assuming all the guesses were correct, for each $i\in[k]$, at the end of the algorithm, we will have selected a center $x_{i}$ such that (i) for each $p\in Y$, if $f^{*}$ assigns $p$ to $i$, then ${\sf d}(p,x_{i})\leq(1+\varepsilon)\cdot{\sf d}(p,o_{i})$, and (ii) capacity of $x_{i}$ is at least that of $o_{i}$, and hence GenBALL-INT is compatible with ASSIGNMENT. This guarantees that there is a feasible assignment from $Y$ to $X$ of cost at most $(1+O(\varepsilon))$ times $\sf OPT$}.
A note on uniform capacities in continuous $\mathbb{R}^{d}$. When capacities are uniform, then the problem can also be defined over continuous Euclidean space, where a facility can be opened at any point in $\mathbb{R}^{d}$. Now, it turns out that the algorithmic scatter dimension of continuous Euclidean space is $O(\tfrac{1}{\varepsilon^{4}}\log(\tfrac{1}{\varepsilon}))$, whereas the non-algorithmic scatter dimension is $1/\varepsilon^{O(d)}$. However, our framework can be adapted to work with algorithmic scatter dimension to yield EPAS for high dimensional continuous Euclidean space as follows. The GenBALL-INT subroutine uses the underlying BALL-INT algorithm admitted by the continuous Euclidean space for every request set $Q_{i},i\in[k]$. However, it could happen that GenBALL-INT selects multiple copies of the same center in its solution $X$, and hence the assignment subroutine (minimum cost max flow subroutine) might fail. To counter this issue, we can adapt the minimum cost max flow subroutine by treating the duplicate centers in $X$ as distinct copies to obtain a feasible assignment, if there exists one. Now, since the algorithmic scatter dimension of this space is $O(\tfrac{1}{\varepsilon^{4}}\log(\tfrac{1}{\varepsilon}))$, the depth of the recursion of our algorithm is also $O(\tfrac{1}{\varepsilon^{4}}\log(\tfrac{1}{\varepsilon}))$, yielding EPAS running time. However, as mentioned before, it could happen that the final solution $X$ returned by our algorithm has multiple copies of the same facility. To handle this, we replace the duplicates with arbitrarily close facilities to the original facility. Note that, such facilities exist since we are in continuous Euclidean space, and it incurs only negligible loss in the approximation factor.
4.1.2 Matroid Clustering
In this variant of $(k,z)$-Clustering, we are additionally given a matroid $\mathbb{M}=(F,\mathcal{I})$, where $\mathcal{I}$ is a collection of independent sets satisfying the matroid axioms181818(1) $\emptyset\in\mathcal{I}$, (2) If $A\in\mathcal{I}$, then for every $B\subseteq A$, it holds that $B\in\mathcal{I}$, and (3) For every $A,B\in\mathcal{I}$ with $|A|<|B|$, there exists some $b\in B\setminus A$ such that $A\cup\left\{b\right\}\in\mathcal{I}$. The goal is to find a subset $C\subseteq F$ of centers that minimizes the $(k,z)$-clustering cost, subject to the constraint that $C$ forms an independent set in the matroid, i.e., $C\in\mathcal{I}$. We assume that the matroid $\mathcal{M}$ is provided via an independence oracle, i.e., given a subset $S\subseteq F$, the oracle reports whether $S\in\mathcal{I}$ in a single step. We assume that the rank of $\mathcal{M}$ is $k$, i.e., we are looking for a subset of facilities that is a base in the matroid $\mathcal{M}$ – this is without loss of generality, since we can always truncate the matroid to rank $k$ if necessary (the independence oracle of the truncated matroid can be easily simulated given the oracle for $\mathcal{M}$). As mentioned earlier, using color coding, we partition $F$ into $k$ colors such that $O$ is colorful. This is nothing but creating a partition matroid $\mathbb{P}$ on $F$. Overall, this means, we are looking for a set of centers in the intersection of $\mathbb{M}$ and $\mathbb{P}$.
The GenBALL-INT algorithm is simply the polynomial time algorithm for the classic problem of matroid intersection [EDM03] of $\mathbb{M}$ and $\mathbb{P}$.
The ASSIGNMENT simply assigns a point completely to the closest center in $X$. In this case, both GenBALL-INT and ASSIGNMENT are compatible with each other.
4.1.3 Fault-tolerant Clustering
We consider (uniform) fault-tolerant setting, which is defined as follows.
Definition 16.
Let $z\geq 1$ be a fixed constant. The input consists of a metric space $\mathcal{M}=(P\cup F,d)$, and positive integers $k\geq 1$ and $\ell\leq k$. The goal is to find $k$ centers $X\subseteq F$ to minimize ${\sf cost}(X):=\sum_{p\in P}{\sf d}_{\ell}(p,X)$, where, for a point $p\in P$, ${\sf d}_{\ell}(p,X):=\min_{X\ \subseteq X,|X^{\prime}|=\ell}\sum_{x^{\prime}\in X^{\prime}}({\sf d}(p,x^{\prime}))^{z}$, is the sum of cost of serving $p$ by $\ell$ nearest centers in $X$.
Now we show how fault-tolerant clustering is capture by our framework. Let $X$ be a set of $k$ centers, and let $X_{p}$ be the set of $\ell$ closest centers in $X$ to $p\in P$. Then, we have that $d_{\ell}(p,X)=\sum_{x\in X_{p}}d(p,x)$, and ${\sf cost}(X)=\sum_{p\in P}\sum_{x\in X_{p}}d(p,x)$. Let $f:P\times X\rightarrow\mathbb{R}_{\geq 0}$ be the corresponding assignment function, i.e., $f(p,x)=1$ if $x\in X_{p}$, and 0 otherwise. In our setting, we will instead work with normalized point-cost $d^{\prime}_{\ell}(p,X)$, which is defined as $d^{\prime}_{\ell}(p,X):=\sum_{x\in X_{p}}\tfrac{d(p,x)}{\ell}$, and the normalized total cost ${\sf cost}^{\prime}(X):=\sum_{p\in P}{\sf d}^{\prime}_{\ell}(p,X)$. Let $f^{\prime}:P\times X\rightarrow\mathbb{R}_{\geq 0}$ be the corresponding assignment function, i.e., $f(p,x)=1/\ell$ if $x\in X_{p}$, and 0 otherwise. Then, note that $\sum_{x\in X}f^{\prime}(p,x)=1$, for every $p\in P$, as required. Moreover,
${\sf cost}(X)=\sum_{p\in P}\sum_{x\in X}f(p,x)(d(p,x))^{z}=\ell\cdot\big(\sum_{p\in P}\sum_{x\in X}f^{\prime}(p,x)(d(p,x))^{z}\big)=\ell\cdot{\sf cost}^{\prime}(X).$
Thus, for any center set $X$, the cost of $f^{\prime}$ is exactly $\ell$ times that of $f^{\prime}$. Moreover, $f^{\prime}$ satisfies the condition of our problem. Hence, in our framework, we instead work with $f^{\prime}$ and the estimate of cost of $(O,f^{\prime})$.
Fault-tolerant $(k,z)$-Clustering is known to admit coreset of size $O(m+(\log n+k+\varepsilon^{-1})k^{2}\varepsilon^{-2z})$ in general metrics due to [HLL+25]. As mentioned earlier, we assume that the facilities are colored in $k$ colors such that $O$ is colorful.
For every request set $Q_{i},i\in[k]$, the GenBALL-INT subroutine returns a feasible facility of color $i$.
Given a set $X$ of $k$ centers, the Assignment subroutine in this case is the following: for a point $p\in Y$ with weight $w(p)$, let $X_{p}\subseteq X$ be the set of $\ell$ closest centers in $X$ to $p$. Then, assign weight $w(p)/\ell$ to each center in $X_{p}$. In this case, both GenBALL-INT and ASSIGNMENT are compatible with each other.
4.2 Unconstrained Center Selection
In this section, we describe some applications of our framework which does not enforce restrictions on the center selection, however, there may be restrictions on how points are assigned to the chosen centers. Therefore, in these applications, our framework yields EPAS for metrics for bounded algorithm $\varepsilon$-scatter dimension, since we can use the best BALL-INT admitted by the metric in our framework.
4.2.1 $(\alpha,\beta)$-Fair Clustering and Variants